OpenStudy (anonymous):

The sum of the angles in a convex 10-sided polygon is: (Points : 2) 1800° 188° 3600° 1440°

OpenStudy (cwrw238):

I always forget the formula for finding this directly i do them using the fact that the sum of the external angles of all convex polygons add up to 360 degrees so here we have 10 angles so ext angle = 360/10 = 36 degrees that means internal angles of regular 10 sided polygon = 180-36 = 144 degrees i leave you to do the rest

OpenStudy (jdoe0001):

\(\bf \textit{sum of internal angles of a polygon}=180({\color{red}{ n}}-2)\qquad {\color{red}{ n}}=\textit{amount of sides}\)